OpenStudy (anonymous):

(-3√24)(5√20) simplify

OpenStudy (lgbasallote):

\[\LARGE (-3\sqrt {24})(5\sqrt{20}) \implies (-3 \times 5) \cdot (\sqrt{24 \times 20})\] does that help?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (lgbasallote):

you're welcome :D